OpenStudy (anonymous):

What value corresponds to the 60th percentile in the following data set? {12, 28, 35, 42, 47, 49, 50} 1) 47 2) 49 3) 28 4) 42

OpenStudy (amistre64):

well, they are in order; what is 60% of 7?

OpenStudy (amistre64):

60/7 * 100 was that it?

OpenStudy (amistre64):

that wasnt it ....

OpenStudy (amistre64):

is it the range we divide up?

OpenStudy (anonymous):

idk thats why im asking

OpenStudy (amistre64):

the formula is on the tip of my brian stem :)

OpenStudy (amistre64):

60/7 is more than 7 so thats bad; is it 7/60?

OpenStudy (amistre64):

nah, thats 11 .....

OpenStudy (amistre64):

60/100 * 7? maybe

OpenStudy (anonymous):

but then how the numbers in the set come into play

OpenStudy (amistre64):

it something simple, that you round up and count that number of points

OpenStudy (amistre64):

ima gonna guess it 47

OpenStudy (anonymous):

nomber 2 is the answer

OpenStudy (amistre64):

if it is #2 then how is that the answer?

OpenStudy (amistre64):

60% of 7 and round up .... sounds right to me

OpenStudy (anonymous):

i think its 47 too total = 263 times .6 brings it to

OpenStudy (amistre64):

7*.6 = 4.2, rounds to 5; #in the 5th spot

OpenStudy (anonymous):

157 - 47 is the closest

OpenStudy (amistre64):

http://en.wikipedia.org/wiki/Percentile

OpenStudy (anonymous):

first of all 60% of 7 is 0.6*7=4.2 meaning is between position 4 & 5 at position 5 u have 47 &position 4 u have 42 formula previous +position(after-previous)*7 =42+0.2(47-42)*7 =49

OpenStudy (anonymous):

let me tell u something in statistics u don't have to say something is to something u have to prove don't guess

OpenStudy (anonymous):

do u understand my answer

OpenStudy (amistre64):

i like your answer better now that I see the math behind it :) but the way i recall it in my stats clretrice comes up to 47 ...

OpenStudy (anonymous):

if u still have questions invite me on ledwaba.reginald@gmail.com

OpenStudy (amistre64):

the wiki would even go: .6*7 = 4.2 4.2 + .5 = 4.7, rnd to nearest is 5 count 5 spots and record the number

OpenStudy (anonymous):

y do u round

OpenStudy (anonymous):

is it 42

OpenStudy (amistre64):

just the way i was taught; you cant pick the 4.2th number so go to the higher one

OpenStudy (amistre64):

its not 42

OpenStudy (amistre64):

42 is the 50th percentile

OpenStudy (anonymous):

so you always round up?

OpenStudy (anonymous):

is not 42 plz dont guess ,don't roud tell the exact value

OpenStudy (amistre64):

yep, by rounding up you include what would have been in the missing spot

OpenStudy (anonymous):

if the value is between 42 and 47 (positions 4 & 5) how can you get 49?

OpenStudy (amistre64):

.5*7 = 3.5; count to 4, 50p = 42

OpenStudy (anonymous):

there r many rational number after 4.2 so u must u the exact number

OpenStudy (amistre64):

@human, i was taught to take the percentage of the data count; and round up, then count to that spot and record the number are there other ways? id say yes :)

OpenStudy (anonymous):

if u studying statistics or actuarial science they r my ways

OpenStudy (amistre64):

your best bet would be to write up, or read up, what the material you have says; so we can interpret it :)

OpenStudy (anonymous):

this is about statistics

OpenStudy (anonymous):

exactly so no guess in ''stats ''

OpenStudy (anonymous):

huh so whats the answer

OpenStudy (amistre64):

this is where a modicum of understanding should come in to allow you to interact in the process in a cogent manner; the answer is what the process gives

OpenStudy (amistre64):

what do you have to bring to the table?

OpenStudy (anonymous):

no disrespect remainder but i'm not convinced its 49 - i'll check my stats nt - i apologise in advance if im wrong

OpenStudy (amistre64):

previous +position(after-previous)*7 =42+ 0.2 (47-42)*7 ^^^ im wondering where the .2 for position comes in; does that reflect the ".2" in the 4.2 ?

OpenStudy (amistre64):

you take the difference between 42 and 47; which is 5 divide it up into 10 even parts and take the 2 of them to get fit it into the spot

OpenStudy (amistre64):

but why do we multiply it by 7 is my wonder ...

OpenStudy (anonymous):

4 is the exact number since we have a remainder u don't ignore it u have to use it . 7 number of values

OpenStudy (amistre64):

yes, 7 data points, but i dont see why we multiply the fitted remainder by 7, much less any other number. ive never seen it in this format before and it makes me curious :)

OpenStudy (amistre64):

this is what i recall :)

OpenStudy (amistre64):

and in examples ...

OpenStudy (amistre64):

My thought process is: 60% of 7 = 4.2 1 2 3 4 5 6 7 12 28 35 42 47 49 50 ^ 4.2th spot; no number so in order to include the 60p, go the the next number on the list 47 is the closest thing we have to an answer then.